3) Yaling decides that she wants to purchase a car in 5 years that will cost her $25,000. She is going to invest money today at a 3% interest rate from her local bank. The bank pays the interest monthly. How much does Yaling need to invest today in order to have $25,000 in five year
Answer #1

We use the formula:
A=P(1+r/12)^12n
where
A=future value
P=present value
r=rate of interest
n=time period.

25000=P*(1+0.03/12)^(12*5)

P=25000/(1+0.03/12)^(12*5)

=25000*0.860869105

=$21521.73(Approx).